# Bidding in Ancillary Service Markets: An Analytical Approach Using Extreme Value Theory

### v2

@@ VII-B Sensitivity analysis
Using this formulation, we can successfully obtain bids in both directions, with $\alpha$ potentially as low as 0.05% for some hours, depending on the underlying distributions. Following the sampling procedure (14) with the current sample set, we would not be able to reduce the constraint violation allowance to less than 6% [22] following the sample-based method, which represents a reduction of only 4%. This highlights the advantage of the analytical reformulation in this type of analysis.
Fig. 7 illustrates the total bids across 24 hours as a function of $\alpha$, as defined in (17), where $\alpha=0.02$ is the lowest value for which we could obtain bids for all 24 hours of the day. The bids shown are averages across 10 runs, similar to the results in Fig. 6. The point-wise 95% confidence level is indicated by a grey band, which is calculated in the standard way using the $t$-distribution [23]:
@@
where $\bar{x}$ and $s$ represent the sample mean and sample standard deviation, respectively, and $t_{.025,9}$ is the critical value from the $t$-distribution with 9 degrees of freedom ($df=10-1=9$) for a two-sided 95% confidence level.
@@
VIII Conclusion and future work
Our findings highlight the effectiveness of the analytical model, derived from concepts of extreme value theory, compared to traditional sample-based methods. While sample-based approaches rely on large datasets to estimate constraints, they often face challenges such as high computational intensity and the potential underestimation of tail risks in rare, extreme market conditions—especially when the failure rate is very small. In contrast, our analytical approach requires fewer samples to reliably estimate the tail of the distribution, ensuring that probability constraints are accurately maintained even in low-probability, high-impact scenarios. This comparison demonstrates that the analytical approach not only reduces computational complexity but also better captures extreme fluctuations, ultimately enhancing both the accuracy and robustness of bids under market volatility. Our case study results illustrate this advantage: using the analytical reformulation, we obtained valid bids with violation rates as low as 0.05%, where the sample-based approach would have required nearly 85,000 samples, according to (14), to successfully place bids with this violation rate. Additionally, the analytical method, incorporating the Bonferroni correction, results in up to 7% fewer constraint violations, providing more reliable bids overall.
In future work, several approaches can further enhance the current study. One promising direction is dependency modeling with copulas to capture correlations between flexibilities, offering deeper insights and optimization strategies. Another opportunity lies in pooling data based on temporal factors, such as distinguishing between weekdays and weekends or accounting for seasonal variations, which could improve model accuracy by reflecting systematic patterns in EV energy consumption. We also suggest exploring dynamic reserve bidding, where strategies adapt in real-time to uncertainties like forecast errors or market fluctuations, to assess their impact on market efficiency and system reliability. Future work could also explore integrating ancillary service price forecasts or surrogate metrics like marginal opportunity costs to further refine the bidding strategy.
@@ Appendix A Flexibility estimation
We assume that the maximum battery capacity for every EV $v\in V$ is the largest energy output exerted by the CB over all charging sessions $s\in S_{v}$, where $S_{v}$ denotes the set of sessions for EV $v$. That is, $L_{v}^{\max}=\max\{\sum_{t}P_{v,t,s}\}$, $t\in[t_{s,0},t_{s}^{\max}]$, $t_{s,0}$ and $t_{s}^{\max}$ being the start and end time of charging session $s$. The SoC at each minute $t$ is given by $L_{v,t}=L_{v}^{\max}-(L_{v,s}-\sum_{i=1}^{t}P_{i,v})$, where $L_{v,s}$ is the total energy applied to the vehicle in the present charging session.
@@
$t_{0}$ and $t_{60}$ being the start and end minute of hour $h$, respectively. Taking the minimum flexibility across the hour in this manner results in a loss of temporal dependencies in the data, which may lead to simplifications in the following analyses. The resulting data contains a year’s worth of flexibility for each hour of the day. Illustrative examples of how the flexibilities are calculated are found in Fig. 1. In Fig. 2 scatter plots of the (minimum) flexibilities for an arbitrary hour over the whole year are shown, where the horizontal line indicates the 10th percentile, which serves as the threshold under which we are selecting data to use for our analysis in the analytical formulation of the bidding strategy.
Acknowledgment
The authors would like to thank Spirii for providing data, as well as Gustav A. Lunde and Emil V. Damm for their extensive work on data preparation, which made the case study possible.

### v4

@@ 7.2 Sensitivity analysis
Using this formulation, we can successfully obtain bids in both directions, with $\alpha$ potentially as low as 0.05% for some hours, depending on the underlying distributions. Following the sampling procedure (14) with the current sample set, we would not be able to reduce the constraint violation allowance to less than 6% (CalafioreCampi) following the sample-based method, which represents a reduction of only 4%. This highlights the advantage of the analytical reformulation in this type of analysis.
Figure 9 illustrates the total bids across 24 hours as a function of $\alpha$, as defined in (17), where $\alpha=0.02$ is the lowest value for which we could obtain bids for all 24 hours of the day. The bids shown are averages across 10 runs, similar to the results in Figure 8. The point-wise 95% confidence level is indicated by a grey band, which is calculated in the standard way using the $t$-distribution (StatBookConfInt):
@@
where $\bar{x}$ and $s$ represent the sample mean and sample standard deviation, respectively, and $t_{.025,9}$ is the critical value from the $t$-distribution with 9 degrees of freedom (df=101=9df=10-1=9) for a two-sided 95% confidence level.
@@
8 Conclusion and future work
This paper has developed an analytical bidding model for an electric vehicle aggregator participating in the Nordic ancillary service markets. The proposed model integrates extreme value theory into a chance-constrained optimization framework, effectively addressing the challenge of meeting Energinet’s P90 requirement while reducing computational burden. Compared with a conventional sample-based benchmark, the approach improves out-of-sample reliability by up to 8% and solves optimization problems up to 4.8 times faster. Furthermore, the methodology successfully generates feasible bids with reliability levels as high as 99.95%, which would otherwise require prohibitively large scenario sets. These findings demonstrate that incorporating extreme value theory can significantly enhance both robustness and tractability in bidding under uncertainty.
The results also highlight several practical considerations for applying the framework in real-world aggregator operations.
  1. 1.
    Data availability: Accurate estimation of flexibility distributions requires high-resolution and reliable data on charging behavior, which may not always be accessible across different fleets.
  2. 2.
    Operational integration: Aggregators must embed the proposed bidding tool into existing decision-making processes that also consider portfolio management, contractual obligations, and financial risk.
  3. 3.
    Implementation and scalability: Real-time application calls for efficient software integration with charging platforms and communication systems to ensure daily bid submissions without excessive computational overhead.
  4. 4.
    Adaptability: Practical deployment may also need to account for temporal correlations in consumption patterns, weekday–weekend variations, and the evolution of charging infrastructure.
This work establishes a regulation-compliant and computationally efficient framework for bidding under uncertainty, offering both methodological innovation and practical insights. Future studies could extend the model by:
  • incorporating price forecasts, activation data, or dynamic reserve bidding;
  • testing the approach within real aggregator platforms; and
  • exploring bidding across multiple markets (e.g., automatic and manual frequency restoration reserves) while considering inter-temporal constraints and dependence structures.
Such developments would further enhance the scalability and applicability of the proposed methodology, supporting reliable and cost-efficient participation of electric vehicles in ancillary service markets and strengthening overall power system security.
@@ Appendix A Flexibility estimation
We assume that the maximum battery capacity for every EV $v\in V$ is the largest energy output exerted by the CB over all charging sessions $s\in S_{v}$, where $S_{v}$ denotes the set of sessions for EV $v$. That is, $L_{v}^{\max}=\max\{\sum_{t}P_{v,t,s}\}$, $t\in[t_{s,0},t_{s}^{\max}]$, $t_{s,0}$ and $t_{s}^{\max}$ being the start and end time of charging session $s$. The SoC at each minute $t$ is given by $L_{v,t}=L_{v}^{\max}-(L_{v,s}-\sum_{i=1}^{t}P_{i,v})$, where $L_{v,s}$ is the total energy applied to the vehicle in the current charging session. By this statement, it is assumed that the SoC at the end of all charging sessions is 100%.
@@
$t_{0}$ and $t_{60}$ being the start and end minute of hour $h$, respectively. Taking the minimum flexibility across the hour in this manner results in a loss of temporal dependencies in the data, which may lead to simplifications in the following analyses. The resulting data contains a year’s worth of flexibility for each hour of the day. Illustrative examples of how the flexibilities are calculated are found in Figure 1. In Figure 2 scatter plots of the (minimum) flexibilities for an arbitrary hour over the whole year are shown, where the horizontal line indicates the 10th percentile, which serves as the threshold under which we are selecting data to use for our analysis in the analytical formulation of the bidding strategy.
Appendix B KS test results for distribution fitting validation
Table 1 provides the mean and standard deviation of the KS test results over 10 runs.
Downward Upward Energy
Hour $D_{n}$ (sd) p-val (sd) $D_{n}$ (sd) p-val (sd) $D_{n}$ (sd) p-val (sd)
1 0.155 (0.023) 0.943 (0.061) 0.118 (0.023) 0.994 (0.005) 0.127 (0.029) 0.981 (0.039)
2 0.195 (0.043) 0.776 (0.193) 0.145 (0.029) 0.956 (0.058) 0.141 (0.026) 0.967 (0.051)
3 0.227 (0.048) 0.621 (0.228) 0.168 (0.031) 0.895 (0.109) 0.182 (0.037) 0.835 (0.149)
4 0.255 (0.038) 0.496 (0.19) 0.109 (0.023) 0.996 (0.005) 0.227 (0.057) 0.627 (0.267)
5 0.273 (0.052) 0.419 (0.252) 0.127 (0.029) 0.981 (0.039) 0.25 (0.058) 0.514 (0.261)
6 0.236 (0.047) 0.591 (0.235) 0.132 (0.026) 0.98 (0.038) 0.205 (0.044) 0.728 (0.193)
7 0.205 (0.039) 0.74 (0.182) 0.123 (0.022) 0.993 (0.005) 0.191 (0.036) 0.8 (0.151)
8 0.195 (0.031) 0.788 (0.139) 0.127 (0.042) 0.958 (0.115) 0.173 (0.019) 0.895 (0.05)
9 0.164 (0.023) 0.919 (0.061) 0.127 (0.042) 0.958 (0.115) 0.155 (0.032) 0.931 (0.116)
10 0.15 (0.037) 0.932 (0.116) 0.136 (0.03) 0.968 (0.051) 0.132 (0.026) 0.98 (0.038)
11 0.145 (0.019) 0.966 (0.05) 0.132 (0.026) 0.98 (0.038) 0.145 (0.047) 0.922 (0.117)
12 0.127 (0.019) 0.992 (0.004) 0.186 (0.026) 0.836 (0.113) 0.127 (0.036) 0.97 (0.052)
13 0.155 (0.032) 0.932 (0.064) 0.123 (0.031) 0.982 (0.039) 0.132 (0.026) 0.98 (0.038)
14 0.127 (0.029) 0.981 (0.039) 0.145 (0.047) 0.921 (0.152) 0.132 (0.026) 0.98 (0.038)
15 0.141 (0.026) 0.967 (0.051) 0.145 (0.019) 0.966 (0.05) 0.136 (0.03) 0.968 (0.051)
16 0.127 (0.029) 0.981 (0.039) 0.141 (0.026) 0.967 (0.051) 0.145 (0.029) 0.956 (0.058)
17 0.141 (0.026) 0.967 (0.051) 0.127 (0.029) 0.981 (0.039) 0.127 (0.029) 0.981 (0.039)
18 0.141 (0.014) 0.978 (0.037) 0.109 (0.023) 0.996 (0.005) 0.109 (0.023) 0.996 (0.005)
19 0.136 (0.03) 0.968 (0.051) 0.136 (0.03) 0.968 (0.051) 0.105 (0.031) 0.986 (0.04)
20 0.127 (0.029) 0.981 (0.039) 0.15 (0.022) 0.955 (0.057) 0.114 (0.024) 0.995 (0.005)
21 0.114 (0.032) 0.984 (0.04) 0.132 (0.026) 0.98 (0.038) 0.132 (0.034) 0.969 (0.052)
22 0.132 (0.026) 0.98 (0.038) 0.141 (0.014) 0.978 (0.037) 0.141 (0.034) 0.957 (0.059)
23 0.173 (0.036) 0.871 (0.138) 0.132 (0.026) 0.98 (0.038) 0.123 (0.031) 0.982 (0.039)
24 0.145 (0.019) 0.966 (0.05) 0.136 (0.0) 0.99 (0.0) 0.132 (0.014) 0.991 (0.003)
Table 1: Mean and standard deviation of the KS test results over 10 runs. The worst values, based on $D_{n}$, are highlighted in red, while the best values are highlighted in blue.